Entry
| Date
| Title
| Site
| Author
| #Graphics
|
203 | Thu 22-Jun-2000 | Sonic summary | | oncley
|
183 | Wed 27-Oct-1999 | Sonic problem at Station3 | stn6 | delany
|
159 | Fri 22-Oct-1999 | Replace ATI At Station 6 | stn6 | jirak
|
129 | Sun 17-Oct-1999 | Sonic put back up at pam 6 | stn6 | semmer
|
128 | Sun 17-Oct-1999 | Change ATI at 20 meters | | semmer
|
123 | Sun 17-Oct-1999 | sonic data at 1.5 and 5 m have interference from hf sheath closures | | miller
|
119 | Sun 17-Oct-1999 | u* from nuw sonics is higher than csat and atik sonics on tower | | horst
|
118 | Sun 17-Oct-1999 | 20m sonic noisy on v axis | tower | horst
|
107 | Sat 16-Oct-1999 | Cycled power on 10m sonic on 55m tower | tower | horst
|
101 | Fri 15-Oct-1999 | Cycled power on 10m sonic on 60m tower | tower | horst
|
98 | Thu 14-Oct-1999 | 20m sonic died; replaced w/ATI (and KH2O from Station 6) | tower | horst
|
86 | Tue 12-Oct-1999 | Tilt of PAM sonics and mean w | stn4 | horst
|
82 | Tue 12-Oct-1999 | ati.55m power cycle | tower | oncley
|
73 | Sun 10-Oct-1999 | 20m ATI was at 10 samp/s | tower | oncley
|
72 | Sun 10-Oct-1999 | replaced 10m ATI | tower | oncley
|
70 | Sun 10-Oct-1999 | 20m boom orientation | tower | horst
|
58 | Fri 08-Oct-1999 | Moving the 40 m boom | tower | lundquist
|
50 | Thu 07-Oct-1999 | ATI sonic gain | | oncley
|
44 | Mon 04-Oct-1999 | Movement of booms | tower | horst
|
28 | Sat 25-Sep-1999 | sonic parameters | tower | oncley
|
- 203: Sonics, Site , Thu 22-Jun-2000 13:51:43 MDT, Sonic summary
The following is pieced together from logbook entries and my memory:
When I arrived, one UW didn't remember rotation coefficients. I reloaded
them. I think we then put this on stn4.
?We also had problems with an ATI which magically fixed itself when I
worked on it?
9/20:
1 NUW5
2 NUW7
3 NUW3
4 ATI3
5 NUW6
6 ATI2
9/25:
10m: 0505
20m: 0202
40m: 0504
6: 0302
spare: NUW4?
1 2 3 4 5 6 10 20 40 spare
NUW5 NUW7 NUW3 0303? NUW6 0302 0505 0202 0504 NUW4
9/27:
- 40m (0504) spiking, removed
- stn4 ATI (3) swapped with spare uw (NUW4?) to have spare ATI (3)
1 2 3 4 5 6 10 20 40 spare
NUW5 NUW7 NUW3 NUW4 NUW6 0302 0505 0202 - 0303
0504s
~9/29:
- 40m replaced with spare () [must have required 3-way swap with 20m due to hotwire mounts?]
1 2 3 4 5 6 10 20 40 spare
NUW5 NUW7 NUW3 NUW4 NUW6 0302 0505 0303 0202 0504s
10/10:
- 10m (0505) spiking, replaced with old 40m (0504)
- 20m () was at 10 samp/s, changed to 20 samp/s
1 2 3 4 5 6 10 20 40 spare
NUW5 NUW7 NUW3 NUW4 NUW6 0302 0504s 0303 0202 0505s
10/14: 20m dead (0303), replaced with sonic from stn6 (0302)
1 2 3 4 5 6 10 20 40 spare
NUW5 NUW7 NUW3 NUW4 NUW6 - 0504s 0302 0202 0505s
0303d
10/15:
- Took bad sonic (0303) back to Boulder,
- Found problem with old 10m (0505) was that it was programmed
like a UW array. Fixed.
1 2 3 4 5 6 10 20 40 spare
NUW5 NUW7 NUW3 NUW4 NUW6 - 0504s 0302 0202 0505
0303[dB]
10/16: blown cap and fuse in 0303? fixed by Gary
1 2 3 4 5 6 10 20 40 spare
NUW5 NUW7 NUW3 NUW4 NUW6 - 0504s 0302 0202 0505
0303[B]
10/17:
- 20m (0302) noisy v, replaced by 0505
- old 20m (0302) with noisy v moved back to stn6
1 2 3 4 5 6 10 20 40 spare
NUW5 NUW7 NUW3 NUW4 NUW6 0302 0504s 0505 0202 0303[B]
?: 0303 returned to field by Cathy & Tony
1 2 3 4 5 6 10 20 40 spare
NUW5 NUW7 NUW3 NUW4 NUW6 0302 0504s 0505 0202 0303
10/22: Replaced stn6 (0302) with spare (0303)
1 2 3 4 5 6 10 20 40 spare
NUW5 NUW7 NUW3 NUW4 NUW6 0303 0504s 0505 0202 0302n
10/27: Stn3 [NUW3?] has bad v&w. No solution found. (Probably bad xducer)
1 2 3 4 5 6 10 20 40 spare
NUW5 NUW7 NUW3b NUW4 NUW6 0303 0504s 0505 0202 0302n
- 183: Sonics, Site stn6, Wed 27-Oct-1999 13:59:30 CDT, Sonic problem at Station3
The UW sonic anemometer v and w were bad for station 3.
At ~1200 all the connectors were checked and the transducers cleaned and
jiggled slightly.
No change was observed.
- 159: Sonics, Site stn6, Fri 22-Oct-1999 11:47:23 CDT, Replace ATI At Station 6
Noted yesterday the V component of the ATI at station 6 had a positive offset compared to rest of the sonics. Presious status has noted peculiarities in the response of this sonic. Replaced with the sensor brought out by TOny and CAthy on the 20th. Old sensor # is ATI 980302 (ATI#2). New Sensor ATI 98303.
- 129: Sonics, Site stn6, Sun 17-Oct-1999 19:28:35 CDT, Sonic put back up at pam 6
Tom installed the sonic back on station 6. The mast was back in its
upright position at 17:15 CDT, Oct 17.
- 128: Sonics, Site , Sun 17-Oct-1999 19:23:54 CDT, Change ATI at 20 meters
Tom and I changed sonics at the 20 meter height. The boom was lowered
at 18:25 CDT, Oct 17.
The sonic at 20 meters is 980505. The sonic that came down went back
to station 6 where it came from a few days ago.
- 123: Sonics, Site , Sun 17-Oct-1999 16:46:37 CDT, sonic data at 1.5 and 5 m have interference from hf sheath closures
The hot film sheath closing every hour for 1 minute can be seen in the variances of the adjacent sonics at 1.5 and 5 m. So the sonic data during the sheath closures is suspect.
- 119: Sonics, Site , Sun 17-Oct-1999 13:32:22 CDT, u* from nuw sonics is higher than csat and atik sonics on tower
With the high winds, u* from the nuw sonics is noticeably higher than
u* for all the sonics on the 55m tower, including those at 1.5, 5 and
10m.
- 118: Sonics, Site tower, Sun 17-Oct-1999 13:03:14 CDT, 20m sonic noisy on v axis
The v axis on the 20m ATIK has been spiking. We plan to move this unit
back to station 6 this afternoon and mount the sonic removed from
10 m on Oct 10. SPO found that this sonic was improperly programmed
as though it were a UW array. He fixed this and we will mount it at
20m.
- 107: Sonics, Site tower, Sat 16-Oct-1999 09:20:04 CDT, Cycled power on 10m sonic on 55m tower
We turned off cosmos temporarily at 1350 CDT, Oct 15, for a test of
the filtering on the 20m KH2O. This stopped the 10m sonic again.
Power was cycled on the 10m sonic at 1715 CDT to restart it.
- 101: Sonics, Site tower, Fri 15-Oct-1999 10:37:09 CDT, Cycled power on 10m sonic on 60m tower
The 10m sonic on the 60m tower quit around 1755 local time, Oct 14.
Restarted it around 1000 local time, Oct 15.
- 98: Sonics, Site tower, Thu 14-Oct-1999 18:04:22 CDT, 20m sonic died; replaced w/ATI (and KH2O from Station 6)
The ATI at 20m died at 10:40 am local time, Oct 14. We removed the
ATI and KH2O from Station 6 at 15:30 local time and installed them at 20m.
The sonic came on line at 16:35.
- 86: Sonics, Site stn4, Tue 12-Oct-1999 15:08:17 CDT, Tilt of PAM sonics and mean w
The mean vertical velocity for PAM station #4 appears anomolous
compared to the other PAM stations. The explanation is apparent
from tilt plots. The sonic at station #4 is tilted about 180
degrees different from those at the other PAM stations.
- 82: Sonics, Site tower, Tue 12-Oct-1999 10:02:48 CDT, ati.55m power cycle
I cycled power on 55m ati which brought it back to life. This done about
9:30 (and I was up the tower to ragwort). This sonic has been down since
2pm yesterday - oops...
- 73: Sonics, Site tower, Sun 10-Oct-1999 17:09:41 CDT, 20m ATI was at 10 samp/s
Somehow, the 20m ATI was set to 10 samp/s. I assume that this has been the
case since set-up. It is now at 20 samp/s.
- 72: Sonics, Site tower, Sun 10-Oct-1999 17:05:40 CDT, replaced 10m ATI
Something has been wrong with the 10m ATI since it was installed. The worst
signal is just w. I just replaced it with our "spare" which was removed from
40m earlier due to spiking and never changed. I'm hoping that this will
work ok in the lower wind speeds at 10m. (It ran on the bench just fine.)
Note that this entailed switching the hotwire mount, which fortunately was
straightforward.
Also note that this sonic also has a level, so the config is the same.
(Unlike when we swapped at 40m -- the sonic now there is reporting dummy
values for the levels (i.e. analog output is turned on, but there is no
signal.))
- 70: Sonics, Site tower, Sun 10-Oct-1999 10:49:33 CDT, 20m boom orientation
I rechecked the 20m boom this morning and found that the mount had
shifted to the south, presumably caused by the process of pushing the
boom to the north when raising it (to avoid crossed guy wires above it).
At 10:20 am local time, Oct 10, I moved the mount to its correct
position that was marked by black electrical tape. Unfortunately,
the boom still appears to be angled a few degrees to the north
relative to the other booms.
- 58: Sonics, Site tower, Fri 08-Oct-1999 16:50:42 CDT, Moving the 40 m boom
This entry should have been made on Tuesday, October 5. From 12:40-13:03 CDT, I raised and lowered the 40 m boom to mount a hot-wire probe next to the sonic.
- 50: Sonics, Site , Thu 07-Oct-1999 16:32:50 CDT, ATI sonic gain
Can change R42 on first stage amplifier. Default value is 33.2K, could
change to 49.9K to obtain higher gain.
- 44: Sonics, Site tower, Mon 04-Oct-1999 10:53:32 CDT, Movement of booms
From 0930 to 1045, Horst and Lundquist raised and lowered the beams at
40m (to remove a hotwire probe shield and support), 20 m (to mount a probe),
and 10 m (to switch out shield/support and mount a probe), in that order.
- 28: Sonics, Site tower, Sat 25-Sep-1999 16:43:03 CDT, sonic parameters
adam chan height type owner serial
cosmos 200 2.5m CSAT Miller
cosmos 201 5m CSAT Miller
cosmos 204 10m ATI NCAR 980505 (w/levels)
cosmos 205 20m ATI NCAR 980202
marigold 200 30m CSAT Mahrt
ragwort 200 40m ATI NCAR 980504 (w/levels)
ragwort 201 50m CSAT Mahrt
ragwort 202 55m ATI PNL 980415
All ATIs:
Sample Options Menu\r\n
A. Median Filter (Off)\r\n
B. Data Quality Algorithm (Off)\r\n
C. Shadow Correction Algorithm (Off)\r\n
D. Remove RH from Temp Calculation (Off)\r\n